OpenStudy (boldjon):

the axis of symmetry for the graph of the function is f(x) = 1/4x^2 + bx + 10 is x = 6. What is the value of b?

OpenStudy (solomonzelman):

axis-of symmetry := the vertical line about which the parabola is symmetric (the vertical line that splits the parabola into 2 equal parts) You need to complete the square (in terms of b), and find the x-coordinate of the vertex (in terms of b), and set this x-coordinate of the vertex equal to 6, to solve for b.

OpenStudy (solomonzelman):

And vertex can be found through «completing the square»

OpenStudy (boldjon):

u want me to plug in the x first?

OpenStudy (solomonzelman):

I want you to complete the square.

OpenStudy (boldjon):

1/16?

OpenStudy (solomonzelman):

EXAMPLE\( : \) If the axis of symmetry for the function \(\color{#000000 }{ \displaystyle f(x)=2x^2+bx+8 }\) is, \(\color{#000000 }{ \displaystyle x=4 }\), then find \(\color{#000000 }{ \displaystyle b }\). SIMPLIFICATION\( : \) The \(\color{#000000 }{ \displaystyle +8 }\) in the \(\color{#000000 }{ \displaystyle f(x) }\) is just a vertical shift (up), therefore, you can tell that the axis of symmetry is going to be the same for \(\color{#000000 }{ \displaystyle f(x)=2x^2+bx+8 }\) and for \(\color{#000000 }{ \displaystyle f(x)=2x^2+bx }\). So, we can just complete the square by, \(\color{#000000 }{ \displaystyle f(x)=2x^2+bx }\) QUESTION\( : \) \(\color{#000000 }{ \displaystyle f(x)=2\left(x^2+\frac{b}{2}x\right) }\) (factored out of 2) Now, you need to find the number that will make your expression inside the parenthesis into a perfect square trinomial (when added into parenthesis). RULE/CONCEPT\( : \) In general, \(\color{#000000 }{ \displaystyle x^2+2ax+a^2=(x+a)^2 }\) is a perfect square trinomial. After comparing, you can tell that in your case, \(\color{#000000 }{ \displaystyle 2a=\frac{b}{2} }\), and therefore, \(\color{#000000 }{ \displaystyle a=\frac{b}{4} \quad \Rightarrow \quad a^2=\frac{b^2}{16} }\) COMPLETING THE SQUARE\( : \) \(\color{#000000 }{ \displaystyle f(x)=2\left(x^2+\frac{b}{2}x\right) }\) \(\color{#000000 }{ \displaystyle f(x)=2\left(x^2+\frac{b}{2}x+\frac{b^2}{16}-\frac{b^2}{16}\right) }\) \(\color{#000000 }{ \displaystyle f(x)=2\left(x^2+\frac{b}{2}x+\frac{b^2}{16}\right)-2\times \frac{b^2}{16} }\) \(\color{#000000 }{ \displaystyle f(x)=2\left(x^2+\frac{b}{2}x+\frac{b^2}{16}\right)-2\times \frac{b^2}{16} }\) Recall that you really have; \(\color{#000000 }{ \displaystyle f(x)=2\left(x^2+2\left(\frac{b}{4}\right)x+\left(\frac{b}{4}\right)^2\right)-2\times \frac{b^2}{16} }\) in a form of general perfect square trinomial, with a=b/4; So you end up with \(\color{#000000 }{ \displaystyle f(x)=2\left(x+\frac{b}{4}\right)^2-2\times \frac{b^2}{16} }\) So your x-axis (in terms of b) is: \(\color{#000000 }{ \displaystyle x{\rm~of~symm:}\quad \quad h=b/4 }\) But, you are also given that the axis of symmetry is, \(\color{#000000 }{ \displaystyle x{\rm~of~symm:}\quad \quad h=4 }\) So, \(\color{#000000 }{ \displaystyle h=b/4,\quad h=4 \quad\quad \Rightarrow \quad4=b/4\quad \Rightarrow \quad \color{blue}{16=b}}\)

OpenStudy (solomonzelman):

This can certainly be written shorter; this isn't a lot, don't be afraid... (I just elaborated everything)

OpenStudy (boldjon):

O_O ok i'll... i'll try

OpenStudy (cmtboy2):

{-_=]

OpenStudy (boldjon):

ohhhhhhh

OpenStudy (boldjon):

thanks

OpenStudy (cmtboy2):

?

OpenStudy (cmtboy2):

https://www.youtube.com/watch?v=WpVChAg2hRg

OpenStudy (cmtboy2):

this i me drag racing

OpenStudy (cmtboy2):

im the red one

OpenStudy (boldjon):

dude, how old are u

OpenStudy (cmtboy2):

12

OpenStudy (boldjon):

ah ok

OpenStudy (boldjon):

ಠ_ಠ

OpenStudy (cmtboy2):

{0___0]

OpenStudy (boldjon):

closing this in 5

OpenStudy (boldjon):

4

OpenStudy (boldjon):

3

OpenStudy (boldjon):

2

OpenStudy (cmtboy2):

ilots that rac the first time then i came back and bett her

OpenStudy (boldjon):

1

OpenStudy (boldjon):

oh ok